# Impulsive mixing of stellar populations in dwarf spheroidal galaxies

### v1

@@ 3.3 Adiabatic evolution of the energy distribution
[…]
Adiabatic invariance of $I$ for radial orbits implies the same scaling of evolved energy $\mathcal{E}$ and dark matter mass $M$ as found for circular orbits using angular momentum conservation (Eq. 37). Hence, the evolved energies coincide for radial and circular orbits in the power-law potential with $\gamma=1$. This is the case of self-similar adiabatic expansion. However, if a stellar tracer expands into a region of differing power-law slope, radial and circular orbits reach new equilibrium configurations that correspond to different energies, causing a slight broadening of the overall energy distribution, as can be seen in the bottom panel of Fig. 7.
Figure 8: Number density profiles $\nu_{\star}(r)$ of stellar tracers with (initial) delta function energy distributions, embedded in a cuspy dark matter halo. Profiles are shown for the initial conditions ($M/M_{0}=1$), as well as for snapshots where the mass sourcing the potential as been adiabatically lowered to $M/M_{0}=1/10$ and $1/100$, respectively. The profiles are normalized by the initial central number density $\nu_{\star 0}(0)$. Left panel: Stellar tracer with initial half-light radius $r_{\mathrm{h0}}/r_{\mathrm{mx0}}=1/64$, deeply embedded within the density cusp of the surrounding halo. The density profile remains well-approximated by Eq. 17 throughout the entire adiabatic evolution, consistent with the energy distribution remaining narrow and the kinematics remaining approximately isotropic. Right panel: Stellar tracer with $r_{\mathrm{h0}}/r_{\mathrm{mx0}}=1/4$. As the stellar tracer expands into a region of steeper power-law slope of the density profile of the underlying halo (see bottom panel of Fig. 2), the inner regions retain isotropic kinematics, whereas at larger radii, the stellar distribution becomes tangentially biased. The same systematics are observed in the case of the cored halo (not shown).
@@ 3.4 Adiabatic evolution of density and anisotropy profile
In contrast, the right panel of Fig. 8 shows the evolved density profiles for a more extended stellar tracer with initial $r_{\mathrm{h0}}/r_{\mathrm{mx0}}=1/4$. This tracer expands into a region with a steeper power-law slope of the underlying potential, and its energy distribution (slightly) broadens. The evolved stellar profile remains centrally isotropic but develops a tangential bias towards the outskirts, which is consistent with (more) circular orbits dominating the least-bound end of the evolved energy distribution (see Fig. 7).
@@ 3.5 Adiabatic evolution of metallicity gradients
Instead, for the case of the metallicity gradient between the extended metal poor population (dark blue symbols in Fig. 5) and the deeply embedded metal rich one (red symbols), $\Delta\mathrm{[Fe/H]}/\Delta r_{\mathrm{h}}$ will drop even more rapidly with decreasing $M/M_{0}$: the more extended (metal poor) population expands in a region of steeper dropping power-law slope of the underlying dark matter halo than the deeply embedded (metal rich) one. Hence, to maintain $r_{\mathrm{h}}M_{\mathrm{h}}=\text{const}$, the extended population expands by a larger factor than the deeply embedded one, causing the gradient between the populations to become even shallower (for a discussion of the formation of shallow metallicity gradients as a consequence of tides, see also Ogiya et al. 2022).
@@
Figure 10: Like Fig. 6 (bottom panel), but showing the relative change in remnant (dark matter) mass $M_{\mathrm{h}}=M(<r_{\mathrm{h}})$ (enclosed within the new equilibrium half-light radius $r_{\mathrm{h}}$ reached after an impulsive change to the underlying potential) as a function of the relative change in half-light radius $r_{\mathrm{h}}/r_{\mathrm{h0}}$. In addition to the data shown in Fig. 9, here we include simulations for even stronger perturbations, spanning $10001\geq M/M_{0}\geq 1/$ 1000. The area of individual points indicates the remnant number of (bound) stars $N_{\star}/N_{\star 0}$. In the linear regime, the final equilibrium states coincide with the adiabatic case. For the non-linear regime, the mass enclosed within the stellar half-light radius increases as the half-light radius expands, until stars are lost.
@@ 4.1 Impulsive evolution of size and kinematics
Analogously to Sec. 3.1, we begin the discussion of the response of stellar systems to impulsive perturbations by measuring the stellar half-light radii $r_{\mathrm{h}}$ and average velocity dispersions $\langle\sigma^{2}\rangle^{1/2}$ in the new equilibrium configuration reached after the impulsive perturbation. We ensure that each stellar population has returned to virial equilibrium (within the limits of numerical resolution of our simulations) when we measure $r_{\mathrm{h}}$ and $\langle\sigma^{2}\rangle^{1/2}$. Fig. 9 shows the results of these experiments. The initial conditions are marked by crossed-out circles along the gray curve labeled $M/M_{0}=1$. All simulations start with these initial conditions. Open circles show the new equilibrium configurations reached after impulsively lowering the (dark matter) mass $M$ sourcing the potential by $-0.2,-0.4,\dots,-2\,\mathrm{dex}$, respectively, in a single step. The left-hand panel of Fig. 9 shows the equilibrium properties for four initially mono-energetic stellar components in a cuspy halo, while the right panel shows the equivalent simulation results for a cored halo. Gray curves show the velocity dispersion expected for a purely mono-energetic stellar component with isotropic kinematics, with the distance between two curves corresponding to a difference of $0.2\,\mathrm{dex}$ in the mass $M$.
@@
Similarly to the adiabatic case, the systematics of the new equilibrium states reached after an impulsive perturbation can be summarized by expressing the amplitude of the evolution of the underlying potential as a fraction of the remnant (dark matter) mass enclosed within the stellar half-light radius, $M_{\mathrm{h}}/M_{\mathrm{h0}}$. Fig. 10 shows the final equilibrium $M_{\mathrm{h}}/M_{\mathrm{h0}}$ as a function of the relative change $r_{\mathrm{h}}/r_{\mathrm{h0}}$ of the stellar half-light radius with respect to its initial value. For $r_{\mathrm{h}}/r_{\mathrm{mx0}}\lesssim 1.5$, the final equilibrium state reached after an impulsive perturbation coincides with the adiabatic case. This regime, labeled “linear” in Fig. 10, corresponds to small changes in the underlying potential. This is consistent with the analytical results discussed in Rozier & Errani (2024), showing that for the case of linear perturbations, the final equilibrium state has no memory of whether the perturbation was impulsive or adiabatic. For larger, “non-linear” changes to the underlying potential, the final equilibrium half-light radius $r_{\mathrm{h}}$ increases together with $M_{\mathrm{h}}/M_{\mathrm{h0}}$ (in contrast to the adiabatic case, where $M_{\mathrm{h}}/M_{\mathrm{h0}}$ decreases as $r_{\mathrm{h}}/r_{\mathrm{h0}}$ increases). As previously, the size of the circles in Fig. 10 scales with the remnant number of (bound) stars $N_{\star}$. For impulsive perturbations that are sufficiently strong to result in the loss of stars, the final half-light radius $r_{\mathrm{h}}$ (computed from the remnant bound stars) saturates at a value corresponding roughly to $r_{\mathrm{h}}\sim r_{\mathrm{mx}}$.
Figure 11: Like Fig. 7, but showing the equilibrium energy distributions of four stellar tracers reached after an impulsive change to the underlying potential. Energies are expressed relative to the initial potential minimum. In contrast to the adiabatic case, here the stellar energy distributions widen considerably when the potential changes. These widening distributions may overlap, energetically mixing the stellar populations. Stars are lost once the stellar energy distributions extend into unbound states (gray area with $\mathcal{E}^{\mathrm{(IC)}}\geq 1$). The bottom panel shows the energy distributions measured in the simulation with $M/M_{0}=1/10$ in units of the evolved potential minimum. Dotted black curves show the energy distributions expected from the conservation of single-particle energies $\mathcal{E}=u^{2}/2+\psi(r)$ in the evolved potential. See Fig. 18 for a version of this plot that includes the case of impulsive deepening of the underlying potential.
@@ 4.2 Impulsive evolution of the energy distribution
In contrast to the adiabatic case, stars are lost for sufficiently large impulsive changes of the underlying potential as the evolved energy distributions may extend beyond the bound/unbound limit of the underlying potential. Fig. 11 shows the energy distributions of four stellar components with initial half-light radii of $r_{\mathrm{h}}/r_{\mathrm{mx0}}=1$, $1/4$, $1/16$ and $1/64$ in dark blue, light blue, orange and red (as shown previously for the adiabatic case in Fig. 7). As before, energies are expressed in units relative to the initial potential minimum, $\mathcal{E}^{\mathrm{(IC)}}=1-E/\Phi_{0}^{\mathrm{(IC)}}$ (see Fig. 18 for a version of this figure with energies normalized to the evolved potential minimum). By construction, the energy distributions are delta functions in the initial conditions (top panel). As the mass sourcing the underlying potential is decreased in a single step to $M/M_{0}=10^{-0.2}$ (second panel from the top), $M/M_{0}=10^{-0.4}$ (third panel) or $M/M_{0}=10^{-1.0}$ (fourth panel), the energy distributions widen, and the energetic spacing between the different components decreases. For sufficiently large impulsive changes to the potential, the initially separated energy distributions may overlap in the evolved state: impulsive changes to the underlying potential can mix stellar components energetically (for an impulsive increase of the mass sourcing the potential, see bottom panels of Fig. 18). The bottom panel of Fig. 11 shows the energy distributions for the simulation with $M/M_{0}=10^{-1.0}$ in units of the evolved potential minimum $\Phi_{0}$. Shown also are the energy distributions computed from the conservation of single-particle energies $\mathcal{E}=u^{2}/2+\psi(r)$ in the evolved potential. This calculation is straightforward at the instant of the perturbation, where the kinetic energy can be computed from the initial potential, $u^{2}/2=\mathcal{E}^{\mathrm{(IC)}}-\psi^{\mathrm{(IC)}}(r)$.
@@
Figure 12: Left-hand panels: number density profile $\nu_{\star}(r)$ (top) and anisotropy $\beta(r)=1-\sigma^{2}_{\mathrm{t}}/(2\sigma^{2}_{\mathrm{r}})$ (bottom) of an initially mono-energetic stellar tracer with $r_{\mathrm{h0}}/r_{\mathrm{mx0}}=1/64$, embedded deeply in the density cusp of a dark matter halo. The (dark matter) mass sourcing the underlying potential is reduced impulsively, in a single step, to values of $\log_{10}M/M_{0}=-0.4$, $-1.0$, $-2.0$. The density profile develops a tail towards larger radii which for $M/M_{0}\lesssim 1/100$ is well-approximated by a power law. The stellar tracer retains centrally-isotropic kinematics, but is heavily radially biased in the outskirts. For the model with $M/M_{0}=1/100$, a slight over-density develops with tangentially-biased kinematics, located close to the initial truncation radius $r_{\mathrm{t0}}$ (Eq. 18). Dotted curves show a model for the outer density profile which assumes purely radial orbits for an energy distribution as in Fig. 11 (see Sec. 4.3 and Appendix B for details). Right-hand panels: same as on the left, but for an extended stellar tracer with $r_{\mathrm{h0}}/r_{\mathrm{mx0}}=1/4$.
@@ 4.3 Impulsive evolution of the density and anisotropy profile
While in the adiabatic case the shape of the stellar density profile evolves only very little, consistent with the stellar components retaining very narrow energy distributions throughout their evolution, in the impulsive case, the stellar density profile evolves considerably in response to an abrupt change of the underlying potential. This is shown in the left panel of Fig. 12, taking as an example a stellar component that initially is deeply embedded within a cuspy dark matter halo ($r_{\mathrm{h0}}/r_{\mathrm{mx0}}=1/64$). The simulations shown here use $N=10^{6}$ particles, and $N=4\times 10^{6}$ in case of the run with $\log_{10}M/M_{0}=-2$. As the mass sourcing the underlying potential is decreased in a single step to $\log_{10}M/M_{0}=-0.4$, $-1.0$, $-2.0$, the density profile of all bound stars develops a tail towards larger radii. For stellar energy distributions that extend all the way to the bound/unbound limit ($\mathcal{E}=1$), the outer (bound) stellar density profile is well approximated by a gradually steepening power law. For heavily perturbed systems and large radii ($r\gg a$), we measure an outer power-law slope of $-4$, consistent with simulation results of violent relaxation (see e.g. Hénon 1964; van Albada 1982 ) and the ejection of stars from globular clusters (Peñarrubia et al., 2017).
@@
While the centers retain isotropic kinematics, the outskirts are radially biased (see bottom panel of Fig. 12). The same systematics are observed for the case of a more extended stellar tracer with an initial half-light radius of $r_{\mathrm{h0}}/r_{\mathrm{mx0}}=1/4$, shown in the right panels of Fig. 12. The energy distribution of this stellar system extends into the unbound region for impulsive changes to the potential of $M/M_{0}=10^{-1}$ and $M/M_{0}=10^{-2}$ (see Fig. 11).
@@
The shape of the outer density profile can be modeled to good accuracy under the assumption of purely radial orbits. The dotted curve in Fig. 12 is calculated by integrating over a distribution of mono-energetic density profiles with purely radial kinematics (see Appendix B), using as weights an energy distribution $\mathrm{d}N_{\star}/\mathrm{d}\mathcal{E}$ as computed in Sec. 4.2 (shown as a dotted curve in Fig. 11). As the assumption of radial orbits breaks down at radii below $r_{\mathrm{t}}$, we plot the model predictions for the density profiles only for radii outside of $r_{\mathrm{t}}$.
@@
For both stellar tracers shown in Fig. 12, for the case of large impulsive changes to the underlying potential, a small over-density emerges at the location of the initial stellar truncation radius $r_{\mathrm{t0}}$. This over-density coincides with a region of tangentially-biased orbits, and consists of particles that were located in the vicinity of $r_{\mathrm{t0}}$, virtually at rest, at the instant of the impulsive perturbations. With (virtually) no kinetic energy at their disposal at the instant of the perturbation, these particles are confined to orbits with apocenters coinciding with $r_{\mathrm{t0}}$. As any remnant radial component of the velocity is exchanged for potential energy at the apocentre in the evolved potential, the remnant family of orbits sourcing the over-density results tangentially biased (see also the phase portrait in Fig. 14, where the caustic seen at at $r_{\mathrm{t0}}$ corresponds to the over-density discussed here).
@@
Figure 13: Time evolution of an initially mono-energetic stellar population with initial half-light radius $r_{\mathrm{h0}}/r_{\mathrm{mx0}}=1/64$, deeply embedded in a cuspy (left) and a cored (right) dark matter halo. The dark matter mass sourcing the potential is lowered impulsively so that both stellar tracers expand by a factor of roughly two with respect to their initial half-light radius. Top row: Evolution of the half-light radius as a function of time. The $25^{\mathrm{th}}$ and $75^{\mathrm{th}}$ percentiles of the stellar radial distributions are enclosed by the gray shaded band. Times are expressed in units of the period $T_{\mathrm{h0}}$ of a circular orbit at the half-light radius at $t=0$. Second row: average velocity dispersion $\langle\sigma^{2}\rangle$. The radial and tangential components are shown separately as gray and yellow curves, respectively. Third row: Virial quantities. total kinetic energy $T$, total potential energy $W$, and their sum $2K+W$. Fourth row: Ratio of mass $M_{\mathrm{h}}$ enclosed within the half-light radius and dynamically estimated mass $M_{\mathrm{est}}=r_{\mathrm{h}}\langle\sigma^{2}\rangle/G$. Bottom row: Half-light radius and velocity dispersion, normalized to their initial values. While the stellar model embedded in the cuspy halo rapidly phase mixes towards an equilibrium configuration, the stellar population embedded in the cored halo undoes several damped oscillations (in energy, size and velocity dispersion) before reaching a phase-mixed equilibrium state. A video version of this figure is available as an arXiv ancillary file.

### v2

@@ 3.3 Adiabatic evolution of the energy distribution
$I_{\gamma=1}\stackrel{{\scriptstyle\begin{subarray}{c}\mathcal{E}\ll 1\end{subarray}}}{{\approx}}\frac{a\sqrt{-\Phi_{0}}}{3\pi}\left(2\mathcal{E}\right)^{3/2}~~~\text{(radial orbit)}.$ (38)
Adiabatic invariance of $I$ for radial orbits implies the same scaling of evolved energy $\mathcal{E}$ and dark matter mass $M$ as found for circular orbits using angular momentum conservation (Eq. 37). Hence, the evolved energies coincide for radial and circular orbits in the power-law potential with $\gamma=1$. This is the case of self-similar adiabatic expansion.
If, however, an initially mono-energetic stellar tracer expands into a region of changing power-law slope of the underlying potential, the energies in the new equilibrium configuration differ between particles on radial and circular orbits. This causes a slight broadening of the overall energy distribution, as shown in the bottom panel of Fig. 7. Coloured histograms show the energy distributions measured in the $N$-body snapshots, while vertical dashed and dotted lines are computed from the conservation of radial action (Eq. 35, for radial robits) and angular momentum (Eq. 27, for circular orbits), respectively. The width of the energy distribution is accurately delimited by these vertical lines.
Figure 8: Number density profiles $\nu_{\star}(r)$ of stellar tracers with (initial) delta function energy distributions, embedded in a cuspy dark matter halo. Profiles are shown for the initial conditions ($M/M_{0}=1$), as well as for snapshots where the mass sourcing the potential as been adiabatically lowered to $M/M_{0}=1/10$ and $1/100$, respectively. The profiles are normalized by the initial central number density $\nu_{\star 0}(0)$. Left panel: Stellar tracer with initial half-light radius $r_{\mathrm{h0}}/r_{\mathrm{mx0}}=1/64$, deeply embedded within the density cusp of the surrounding halo. The density profile remains well-approximated by Eq. 17 throughout the entire adiabatic evolution, consistent with the energy distribution remaining narrow and the kinematics remaining approximately isotropic. Right panel: Stellar tracer with $r_{\mathrm{h0}}/r_{\mathrm{mx0}}=1/4$. As the stellar tracer expands into a region of steeper power-law slope of the density profile of the underlying halo (see bottom panel of Fig. 2), the inner regions retain isotropic kinematics, whereas at larger radii, the stellar velocity distribution becomes tangentially biased. The same systematics are observed in the case of the cored halo (not shown).
@@ 3.4 Adiabatic evolution of density and anisotropy profile
In contrast, the right panel of Fig. 8 shows the evolved density profiles for a more extended stellar tracer with initial $r_{\mathrm{h0}}/r_{\mathrm{mx0}}=1/4$. This tracer expands into a region with a steeper power-law slope of the underlying potential, and its energy distribution (slightly) broadens. The evolved stellar profile remains centrally isotropic but develops a tangential bias toward the outskirts. Some intuition on the origin of the tangential anisotropy can be gained from the energy distribution in the new equilibrium state. Circular orbits dominate the least-bound end of the evolved energy distribution (see Fig. 7), whereas radial orbits delimit the most-bound end of the energy distribution, which is a consequence of the conservation of angular momentum and radial action, respectively (see Sec. 3.3 for details). The overabundance of circular orbits at the least-bound energies manifests itself as a tangential bias in the velocity dispersion at large radii.
@@ 3.5 Adiabatic evolution of metallicity gradients
Instead, for the case of the metallicity gradient between the extended metal poor population (dark blue symbols in Fig. 5) and the deeply embedded metal rich one (red symbols), $|\Delta\mathrm{[Fe/H]}/\Delta r_{\mathrm{h}}|$ will decrease even more rapidly with decreasing $M/M_{0}$: the more extended (metal poor) population expands in a region of steeper dropping power-law slope of the underlying dark matter halo than the deeply embedded (metal rich) one. Hence, to maintain $r_{\mathrm{h}}M_{\mathrm{h}}=\text{const}$, the extended population expands by a larger factor than the deeply embedded one, causing the gradient between the populations to become even shallower (for a discussion of the formation of shallow metallicity gradients as a consequence of tides, see also Ogiya et al. 2022).
@@
Figure 10: Like Fig. 6 (bottom panel), but showing the relative change in remnant (dark matter) mass $M_{\mathrm{h}}=M(<r_{\mathrm{h}})$ (enclosed within the new equilibrium half-light radius $r_{\mathrm{h}}$ reached after an impulsive change to the underlying potential) as a function of the relative change in half-light radius $r_{\mathrm{h}}/r_{\mathrm{h0}}$. In addition to the data shown in Fig. 9, here we include simulations for even stronger perturbations, spanning $1\geq M/M_{0}\geq 1/1000$. The area of individual points indicates the remnant number of (bound) stars $N_{\star}/N_{\star 0}$. In the linear regime, the final equilibrium states coincide with the adiabatic case. For the nonlinear regime, the mass enclosed within the stellar half-light radius increases as the half-light radius expands, until stars are lost.
@@ 4.1 Impulsive evolution of size and kinematics
Analogously to Sec. 3.1, we begin the discussion of the response of stellar systems to impulsive perturbations by measuring the stellar half-light radii $r_{\mathrm{h}}$ and average velocity dispersions $\langle\sigma^{2}\rangle^{1/2}$ in the new equilibrium configuration reached after the impulsive perturbation. We ensure that each stellar population has returned to virial equilibrium (within the limits of numerical resolution of our simulations) when we measure $r_{\mathrm{h}}$ and $\langle\sigma^{2}\rangle^{1/2}$. Fig. 9 shows the results of these experiments. The initial conditions are marked by crossed-out circles along the gray curve labeled $M/M_{0}=1$. All simulations start with these initial conditions. Open circles show the new equilibrium configurations reached after impulsively lowering the (dark matter) mass $M$ sourcing the potential by $0.2,-0.4,\dots,-2\,\mathrm{dex}$, respectively, in a single step. The left-hand panel of Fig. 9 shows the equilibrium properties for four initially mono-energetic stellar components in a cuspy halo, while the right panel shows the equivalent simulation results for a cored halo. Gray curves show the velocity dispersion expected for a purely mono-energetic stellar component with isotropic kinematics, with the distance between two curves corresponding to a difference of $0.2\,\mathrm{dex}$ in the mass $M$.
@@
Similarly to the adiabatic case, the systematics of the new equilibrium states reached after an impulsive perturbation can be summarized by expressing the amplitude of the evolution of the underlying potential as a fraction of the remnant (dark matter) mass enclosed within the stellar half-light radius, $M_{\mathrm{h}}/M_{\mathrm{h0}}$. Fig. 10 shows the final equilibrium $M_{\mathrm{h}}/M_{\mathrm{h0}}$ as a function of the relative change $r_{\mathrm{h}}/r_{\mathrm{h0}}$ of the stellar half-light radius with respect to its initial value. For $r_{\mathrm{h}}/r_{\mathrm{mx0}}\lesssim 1.5$, the final equilibrium state reached after an impulsive perturbation coincides with the adiabatic case. This regime, labeled “linear” in Fig. 10, corresponds to small changes in the underlying potential. This is consistent with the analytical results discussed in Rozier & Errani (2024), showing that for the case of linear perturbations, the final equilibrium state has no memory of whether the perturbation was impulsive or adiabatic. For larger, “nonlinear” changes to the underlying potential, the final equilibrium half-light radius $r_{\mathrm{h}}$ increases together with $M_{\mathrm{h}}/M_{\mathrm{h0}}$ (in contrast to the adiabatic case, where $M_{\mathrm{h}}/M_{\mathrm{h0}}$ decreases as $r_{\mathrm{h}}/r_{\mathrm{h0}}$ increases). As previously, the size of the circles in Fig. 10 scales with the remnant number of (bound) stars $N_{\star}$. For impulsive perturbations that are sufficiently strong to result in the loss of stars, the final half-light radius $r_{\mathrm{h}}$ (computed from the remnant bound stars) saturates at a value corresponding roughly to $r_{\mathrm{h}}\sim r_{\mathrm{mx}}$.
Figure 11: Like Fig. 7, but showing the equilibrium energy distributions of four stellar tracers reached after an impulsive change to the underlying potential. Energies are expressed relative to the initial potential minimum. In contrast to the adiabatic case, here the stellar energy distributions widen considerably when the potential changes. These widening distributions may overlap, energetically mixing the stellar populations. Stars are lost once the stellar energy distributions extend into unbound states (gray area with $\mathcal{E}^{\mathrm{(IC)}}\geq 1$). The bottom panel shows the energy distributions measured in the simulation with $M/M_{0}=1/10$ in units of the evolved potential minimum. Dotted black curves show the energy distributions expected from the conservation of single-particle energies $\mathcal{E}=u^{2}/2+\psi(r)$ in the evolved potential (see Eq. 40). See Fig. 18 for a version of this plot that includes the case of impulsive deepening of the underlying potential.
@@ 4.2 Impulsive evolution of the energy distribution
In contrast to the adiabatic case, stars are lost for sufficiently large impulsive changes of the underlying potential as the evolved energy distributions may extend beyond the bound/unbound limit of the underlying potential. Fig. 11 shows the energy distributions of four stellar components with initial half-light radii of $r_{\mathrm{h}}/r_{\mathrm{mx0}}=1$, $1/4$, $1/16$ and $1/64$ in dark blue, light blue, orange and red (as shown previously for the adiabatic case in Fig. 7). As before, energies are expressed in units relative to the initial potential minimum, $\mathcal{E}^{\mathrm{(IC)}}=1-E/\Phi_{0}^{\mathrm{(IC)}}$ (see Fig. 18 for a version of this figure with energies normalized to the evolved potential minimum). By construction, the energy distributions are delta functions in the initial conditions (top panel). As the mass sourcing the underlying potential is decreased in a single step to $M/M_{0}=10^{-0.2}$ (second panel from the top), $M/M_{0}=10^{-0.4}$ (third panel) or $M/M_{0}=10^{-1.0}$ (fourth panel), the energy distributions widen, and the energetic spacing between the different components decreases. For sufficiently large impulsive changes to the potential, the initially separated energy distributions may overlap in the evolved state: impulsive changes to the underlying potential can mix stellar components energetically (for an impulsive increase of the mass sourcing the potential, see bottom panels of Fig. 18).
The bottom panel of Fig. 11 shows the energy distributions for the simulation with $M/M_{0}=10^{-1.0}$ in units of the evolved potential minimum $\Phi_{0}$. Shown also are the energy distributions computed from the conservation of single-particle energies $\mathcal{E}=u^{2}/2+\psi(r)$ in the evolved potential. Stars are treated as mass-less tracers of the underlying potential, hence this calculation is straightforward at the instant of the perturbation, where the kinetic energy can be computed from the initial potential, $u^{2}/2=\left[\Phi_{0}^{\mathrm{(IC)}}/\Phi_{0}\right]~\left[\mathcal{E}_{\star}^{\mathrm{(IC)}}-\psi^{\mathrm{(IC)}}(r)\right]$. Note that while in the initial conditions the stars have a mono-energetic energy distribution, the distribution broadens in response to the impulsive perturbation: the change in potential energy of each star depends on it’s position $r$ at the instant of the perturbation. For the type of impulsive perturbations studied here, the perturbed energy distribution then simply reads
$\frac{\mathrm{d}N_{\star}}{\mathrm{d}\mathcal{E}}=\frac{\mathrm{d}N_{\star}}{\mathrm{d}r}~\left|\frac{\mathrm{d}\mathcal{E}}{\mathrm{d}r}\right|^{-1}$ (40)
where $\mathrm{d}N_{\star}/\mathrm{d}r=4\pi r^{2}\nu_{\star}(r)$ follows from the (initial, mono-energetic) stellar density profile, and $|\mathrm{d}\mathcal{E}/\mathrm{d}r|>0$ is the Jacobian of the mapping between a star’s position $r$ at the instant of the perturbation and it’s energy $\mathcal{E}$ after the perturbation (for the type of perturbations studied here, the Jacobian does not vanish).
@@
Figure 12: Left-hand panels: number density profile $\nu_{\star}(r)$ (top) and anisotropy $\beta(r)=1-\sigma^{2}_{\mathrm{t}}/(2\sigma^{2}_{\mathrm{r}})$ (bottom) of an initially mono-energetic stellar tracer with $r_{\mathrm{h0}}/r_{\mathrm{mx0}}=1/64$, embedded deeply in the density cusp of a dark matter halo. The (dark matter) mass sourcing the underlying potential is reduced impulsively, in a single step, to values of $\log_{10}M/M_{0}=-0.4$, 1.0-1.0, 2.0-2.0. The density profile develops a tail toward larger radii which for $M/M_{0}\lesssim 1/100$ is well-approximated by a power law. The stellar tracer retains centrally-isotropic kinematics, but is heavily radially biased in the outskirts. For the model with $M/M_{0}=1/100$, a slight overdensity develops with tangentially-biased kinematics, located close to the initial truncation radius $r_{\mathrm{t0}}$ (Eq. 18). Dotted curves show a model (Eq. B5) for the outer density profile that assumes purely radial orbits for an energy distribution as in Fig. 11 (see Sec. 4.3 and Appendix B for details). Right-hand panels: same as on the left, but for an extended stellar tracer with $r_{\mathrm{h0}}/r_{\mathrm{mx0}}=1/4$. At large radii, the log-slope of the stellar density profile approaches $\mathrm{d}\ln\nu_{\star}(r)/\mathrm{d}\ln r\rightarrow-4$.
@@ 4.3 Impulsive evolution of the density and anisotropy profile
While in the adiabatic case the shape of the stellar density profile evolves only very little, consistent with the stellar components retaining very narrow energy distributions throughout their evolution, in the impulsive case, the stellar density profile evolves considerably in response to an abrupt change of the underlying potential. This is shown in the left panel of Fig. 12, taking as an example a stellar component that initially is deeply embedded within a cuspy dark matter halo ($r_{\mathrm{h0}}/r_{\mathrm{mx0}}=1/64$). The simulations shown here use $N=10^{6}$ particles, and $N=4\times 10^{6}$ in case of the run with $\log_{10}M/M_{0}=-2$. As the mass sourcing the underlying potential is decreased in a single step to $\log_{10}M/M_{0}=-0.4$, 1.0-1.0, 2.0-2.0, the density profile of all bound stars develops a tail toward larger radii. For stellar energy distributions that extend all the way to the bound/unbound limit ($\mathcal{E}=1$), the outer (bound) stellar density profile is well approximated by a gradually steepening power law. For heavily perturbed systems and large radii ($r\gg a$), we measure an outer power-law slope of $\mathrm{d}\ln\nu_{\star}(r)/\mathrm{d}\ln r\approx-4$, consistent with simulation results of violent relaxation (see e.g. Hénon 1964; van Albada 1982 ) and the ejection of stars from globular clusters (Peñarrubia et al., 2017).
@@
Whereas the centers retain isotropic kinematics, the outskirts are radially biased (see bottom panel of Fig. 12). The origin of the radial bias in the outskirts is intuitively explained by noting that the impulsive perturbation studied here is spherical: single-star angular momenta are conserved and limited by the (isotropic) initial conditions. Stars that in response to the impulsive perturbation find themselves on weakly bound orbits that can reach large radii therefore have large orbital eccentricities and limited tangential velocity. The same systematics are observed for the case of a more extended stellar tracer with an initial half-light radius of $r_{\mathrm{h0}}/r_{\mathrm{mx0}}=1/4$, shown in the right panels of Fig. 12. Note that the energy distribution of this stellar system extends into the unbound region for impulsive changes to the potential of $M/M_{0}=10^{-1}$ and $M/M_{0}=10^{-2}$ (see Fig. 11).
@@
The shape of the outer density profile can be modeled to good accuracy under the assumption of purely radial orbits. The dotted curve in Fig. 12 is calculated by integrating over a distribution of mono-energetic density profiles with purely radial kinematics (Eq. B5), using as weights an energy distribution $\mathrm{d}N_{\star}/\mathrm{d}\mathcal{E}$ as computed in Sec. 4.2 (Eq. 40, shown as a dotted curve in Fig. 11). See Appendix B for details. For stellar energy distributions that extend all the way to the bound/unbound limit, at large radii, the log-slope of the model density profile steepens to $\mathrm{d}\ln\nu_{\star}(r)/\mathrm{d}\ln r\rightarrow-4$ (see Eq. B6), consistent with the $N$-body results. As the assumption of radial orbits breaks down at radii below $r_{\mathrm{t}}$, we plot the model predictions for the density profiles only for radii outside of $r_{\mathrm{t}}$.
@@
For both stellar tracers shown in Fig. 12, in the case of large impulsive changes to the underlying potential, a small overdensity emerges at the location of the initial stellar truncation radius $r_{\mathrm{t0}}$. This overdensity coincides with a region of tangentially-biased orbits, and consists of particles that were located in the vicinity of $r_{\mathrm{t0}}$, virtually at rest, at the instant of the impulsive perturbations. With (virtually) no kinetic energy at their disposal at the instant of the perturbation, these particles are confined to orbits with apocenters coinciding with $r_{\mathrm{t0}}$. As any remnant radial component of the velocity is exchanged for potential energy at the apocentre in the evolved potential, the remnant family of orbits sourcing the overdensity results tangentially biased (see also the phase portrait in Fig. 14, where the caustic seen at at $r_{\mathrm{t0}}$ corresponds to the overdensity discussed here).
@@
Figure 13: Time evolution of an initially mono-energetic stellar population with initial half-light radius $r_{\mathrm{h0}}/r_{\mathrm{mx0}}=1/64$, deeply embedded in a cuspy (left) and a cored (right) dark matter halo. The dark matter mass sourcing the potential is lowered impulsively so that both stellar tracers expand by a factor of roughly two with respect to their initial half-light radius. Top row: Evolution of the half-light radius as a function of time. The $25^{\mathrm{th}}$ and $75^{\mathrm{th}}$ percentiles of the stellar radial distributions are enclosed by the gray shaded band. Times are expressed in units of the period $T_{\mathrm{h0}}$ of a circular orbit at the half-light radius at $t=0$. Second row: average velocity dispersion $\langle\sigma^{2}\rangle$. The radial and tangential components are shown separately as gray and yellow curves, respectively. Third row: Virial quantities. total kinetic energy $T$, total potential energy $W$, and their sum $2K+W$. Fourth row: Ratio of mass $M_{\mathrm{h}}$ enclosed within the half-light radius and dynamically estimated mass $M_{\mathrm{est}}=r_{\mathrm{h}}\langle\sigma^{2}\rangle/G$. Bottom row: Half-light radius and velocity dispersion, normalized to their initial values. While the stellar model embedded in the cuspy halo rapidly phase mixes toward an equilibrium configuration, the stellar population embedded in the cored halo undoes several damped oscillations (in energy, size and velocity dispersion) before reaching a phase-mixed equilibrium state. For videos supplementing this figure see Appendix C.